# Soft information decoding with superconducting qubits

### v1

@@ I Introduction
Quantum error correction (QEC) is fundamentally recognized as a central strategy for deploying fault-tolerant computations at scale. It promises exponential error suppression, closing the divide between attainable error rates and those required for practical quantum computing. In recent years, significant progress has been achieved in quantum error correction experiments, demonstrating its emerging capabilitiesKrinner et al. (2022); et al. (2022a); Google Quantum AI (2023); Bluvstein et al. (2024); da Silva et al. (2024); Gupta et al. (2024); Acharya et al. (2024).
These advancements are particularly evident in implementing the surface code, a leading candidate for fault-tolerant quantum computation. Progress has been marked by an upward path from implementing a distance-three ($d=3$) surface code Krinner et al. (2022); et al. (2022a), advancing to a distance-five ($d=5$) surface code Google Quantum AI (2023), and culminating in a distance-seven ($d=7$) surface code that has also shown evidence for improvements in the fidelities of logical entangling gates with increasing code distance Bluvstein et al. (2024). Furthermore, recent experiments claim to have achieved break-even error correction performance, matching or surpassing their physical counterparts. Key milestones include achieving break-even fidelity in the initialization of logical qubits within a color code Bluvstein et al. (2024), preparing logical Bell pairs with error rates 800 times lower than their physical level da Silva et al. (2024), and successfully distilling magic states with fidelity beyond the break-even point Gupta et al. (2024). Furthermore, in recent experiments, below surface threshold performance was achieved on a distance-seven ($d=7$) code Acharya et al. (2024). These claims highlight quantum error correction’s growing effectiveness and practical relevance.
@@
Pattison et al. developed a method to circumvent this loss by incorporating the full range of analog data, also known as soft information, into the decoding process Pattison et al. (2021). Their simulations indicated that using soft information in surface code decoding can significantly raise the code’s threshold compared to decoders that utilize only binary, or “hard”, information. This result reveals a substantial improvement in the exponential scaling of error suppression, going beyond just a constant benefit. The findings underscore the importance of utilizing all available information from the experimental setup, thus optimizing quantum error correction strategies.
In addition to the work by Pattison et al., the benefit of employing soft information has been validated through simulations using neural network decoders Varbanov et al. (2023); Bausch et al. (2023). On the hardware front, soft decoding has been implemented in different settings: for improving the readout fidelity on a Si/SiGe two-qubit system Xue et al. (2020), and for superconducting qubits using both a distance-three ($d=3$) heavy-hexagon code Sundaresan et al. (2023) and, recently, a distance-three ($d=3$) two-dimensional repetition code Ali et al. (2024). However, these hardware experiments were limited to at most $d=3$ codes, which restricted their ability to evaluate and compare thresholds between soft and hard decoding—a significant and anticipated aspect of this research.
Furthermore, the experiments of Refs. Sundaresan et al. (2023); Ali et al. (2024), involving $T=10$ and 16 rounds of stabilizer measurements, respectively, managed to post-select for leakage errors—transitions to higher energy states detrimental to quantum error correction et al. (2022b). This post-selection was feasible because the low number of measurement rounds helped restrict the accumulation of leakage, which typically increases with the duration of the experiment. Consequently, the issue of leakage and its impact on soft decoding, also not addressed in the initial studies by Pattison et al. or subsequent simulation efforts, remains a critical area for further investigation.
@@ II.1 Repetition codes
After measuring the stabilizer, one might consider resetting the ancilla to the $|0\rangle$ state to prevent additional error propagation. However, resetting the qubit is a non-trivial task, requiring either waiting for the qubit to decay to the ground state or applying an active reset operation. The former would mean that the code qubits in the system must be idle for as long as the ancillas decay, causing them to decay themselves. The latter would require applying an $X$ gate conditioned on the measurement outcome. However, this would require dynamic operations, which also cause idling times and specific errors. To circumvent these issues, we do not reset the ancilla qubits. Moreover, it has been shown that omitting ancilla resets can have benefits in the decoding of memory experiments Gehér et al. (2024), further justifying our approach.
@@ II.2 Decoding
Repetition codes are almost cyclic, enabling their decoding through the minimum weight perfect matching (MWPM) algorithm Higgott (2021); Higgott and Gidney (2023). This decoding method is also extensively used for surface codes, widely recognized for their promising potential in fault-tolerant quantum error correction Krinner et al. (2022); et al. (2022a); Google Quantum AI (2023); Bluvstein et al. (2024); Acharya et al. (2024). Therefore, improved decoding performance achieved with repetition codes likely has implications for the broader application with surface codes.
@@
Error chains, i.e., a set of errors with overlapping syndromes, will induce either one syndrome—if the chain terminates at a boundary—or two syndromes. Therefore the decoding problem is equivalent to finding the shortest path on the graph to pair up all the nontrivial syndromes. The matching decoder solves this problem on a weighted graph, where the edge weights $w=\log((1-p)/p)$ are calculated using the probability $p$ of the corresponding error event. The minimum-weight set of edges found by MWPM then corresponds to the most likely set of errors Higgott (2021); Higgott and Gidney (2023).
We typically evaluate the merit of a quantum code based on its threshold value $p_{th}$, defined as the upper limit of error probability at which the code can still suppress errors. This value originates in the threshold theorem Nielsen and Chuang (2010), which posits that assuming a constant probability of a fault $p$ across all potential errors, the logical error rate $\epsilon_{L}$ decreases exponentially with respect to the order of fault-tolerance $f$, which is the maximum number of faults that the code can correct Ghosh et al. (2012). For repetition codes with distance $d$, the order of fault-tolerance is $f=\lfloor d/2\rfloor$ as the code can detect up to $d-1$ errors but only unambiguously correct up to $\lfloor d/2\rfloor$ errors. Consequently, the logical error rate $\epsilon_{L}$ exponentially decreases with the distance of the code as
@@
The code in question does not solely determine the threshold value, as it is also significantly influenced by the decoding method Google Quantum AI (2021). Typically, researchers simulate the performance of a code and its decoder across varying error rates to estimate the threshold. Moreover, the error parameters in the circuit-level error model do not need to be equal, requiring a threshold parameter for each error mechanism Hetényi and Wootton (2024). Determining these thresholds in experiments would require precise manipulation of the noise model and, hence, gates with accurately controllable error rates—a daunting task.
Instead, we employ the error suppression factor $\Lambda$ as a practical metric for comparing the performance of various codes and decoders et al. (2015). We define $\Lambda$ through the scaling of the logical error rate $\epsilon_{L}$ as
@@
When decoding multiple rounds of syndrome measurements, we obtain the total logical error rate $P_{L}$ rather than the previously discussed logical error rate per round $\epsilon_{L}$. Assuming each round incurs a logical error with probability $\epsilon_{L}$, the total logical error rate $P_{L}$ after $T$ rounds is given by Google Quantum AI (2021):
@@ II.3 Soft information
In superconducting hardware, we perform qubit readouts by coupling a resonator to the qubit. The qubit’s state then induces a dispersive shift in the resonator’s frequency. To detect this shift, we apply a microwave tone to the resonator and measure the reflected signal, which has an altered amplitude and phase according to the qubit’s state. We integrate this reflected signal to extract two key metrics: the in-phase ($I$) and quadrature ($Q$) components and plot them on the IQ plane Bronn et al. (2017). This plane has $I$ and $Q$ on its axes and visually represents the qubit’s state through the distribution of $I$ and $Q$ points. Statistically, the $I$ and $Q$ points cluster into two distinct Gaussian distributions on the IQ plane, each representing one of the qubit’s possible states. Fig. 2 shows statistical IQ data obtained from a superconducting qubit. To ascertain the qubit’s state from the reflected signal, we assess the likelihood of a data point belonging to each Gaussian and convert it into a binary outcome. The dispersive shift determines the separation between the two Gaussians, whereas the width of the Gaussian curves is dependent, among other things, on the duration of the integration window. Despite efforts to separate these distributions, a non-zero overlap exists, creating a situation where an IQ data point might fall into an area more likely associated with the incorrect state. Such occurrences lead to a misassignment error, commonly called a soft flip.
@@
This assumption generally holds for stabilizer measurements after sufficient rounds, as the ancilla qubits will likely be randomly in the $|0\rangle$ or $|1\rangle$ state. However, for the initial rounds and particularly for the final logical code readout, this assumption may not hold. Therefore, incorporating additional information derived from Pauli tracing Gidney (2021) could improve the model’s accuracy and, consequently, the efficacy of the methods discussed in this chapter.
@@ II.6 Soft decoding in the presence of leakage
Transmon qubits, featured in IBM’s devices, are a leading platform for quantum computation. However, transmons have a finite anharmonicity, allowing transitions to higher energy states. Such transitions can excite a qubit, resulting in leakage errors. These excitations often occur when implementing two-qubit gates Motzoi et al. (2009); Chen et al. (2016), as well as measurements Barends et al. (2014); Yan et al. (2018); Rol et al. (2019); Negîrneac et al. (2021).
Leakage poses a significant challenge in quantum error correction. Once a qubit leaks, it does not operate as intended, leading to persistent faults over multiple stabilizer measurements. Quantum error correction typically assumes that errors are uncorrelated; however, leakage undermines this assumption, posing a significant challenge for the experimental implementation of quantum error correction protocols. Miao et al. demonstrated the detrimental effect of leakage on decoding performance in superconducting qubits et al. (2022b), showing a much higher sensitivity of code performance to leakage than Pauli errors. Past experiments involving multiple rounds of quantum error correction have focused on identifying and post-selecting leaked qubits, as seen in the studies conducted by Google Quantum AI (2021); Krinner et al. (2022).
Soft information decoding experiments have also selectively removed samples with detected leaked qubits Ali et al. (2024); Sundaresan et al. (2023). Notably, these studies involved no more than $T=16$ rounds of measurements. In contrast, our experiments extend to $T=100$ rounds, significantly increasing the likelihood of leakage and thus reducing the yield to nearly zero if post-selection were applied.
Several strategies have been developed to mitigate leakage in quantum systems, such as implementing leakage-reducing reset protocols Google Quantum AI (2021); et al. (2021), applying optimal control techniques to gate operations Werninghaus et al. (2021), and integrating leakage reduction units directly into the circuit design Suchara et al. (2014); Fowler (2013). In line with our decision to use IBM Quantum’s deployed device IBM Sherbrooke, which is not optimized for quantum error correction, we explored the extreme case where no leakage mitigation is applied at the hardware level. Instead of focusing on implementing sophisticated leakage-reduction protocols, we investigated how effectively leakage can be managed solely through classical decoding techniques. As Fowler et al. Fowler (2013) showed, a threshold still exists under leakage, although significantly lower, making this a compelling test of soft decoding’s robustness under realistic, unmitigated conditions.
Leakage manifests as a separate Gaussian distribution in the IQ plane, typically positioned between the Gaussians representing the computational states Sundaresan et al. (2023). This separation allows us to detect leakage using the IQ data already available, enabling the classification of leaked qubits based on their IQ position.
@@
Second, as described in Sundaresan et al. (2023), the Gaussian distribution of the leaked state has a variance similar to the computational states, ideally calling for a linear decision boundary between leaked and computational states. Our approach, based on outlier detection, results in a circular decision boundary, which tends to underestimate the number of leaked states, as shown in Fig. 6. Addressing this limitation would ideally involve including a distinct preparation of leaked states in the calibration data and performing a three-state discrimination process, as described in Sundaresan et al. (2023).
@@ III.1 Simulations
To explore the advantages of soft decoding repetition codes on superconducting hardware, we initially performed simulations that mimic the behavior of realistic quantum hardware to assess the effectiveness of soft decoding. After these preliminary simulations, we conducted experiments on IBM Quantum’s deployed device IBM Sherbrooke, a 127-qubit Eagle device not specifically tuned for quantum error correction tasks like repetition codes. Rather than mitigate leakage and other measurement errors through fine-tuning of the measurement process, as would be done in a full demonstration of an error correcting code, we instead simply use the measurements as provided “out-of-the-box” through Qiskit Javadi-Abhari et al. (2024). This allows us to investigate how much the leakage can be mitigated purely through the decoding. Using this general-purpose approach, we did not aim to maximize QEC performance but to evaluate how soft decoding behaves under practical conditions comparing the outcomes with those of the simulations.
@@
While this setup with $T=50$ does not approach the condition of $T\gg d$, it aligns with the conventional experimental practice where $d$ is set equal to $T$, ensuring that the “time distance” equals the code distance. Ref. Google Quantum AI (2021) showed that time boundary effects in repetition codes do not affect the logical error rate per round for $T\geq 10$, supporting our decision to set $T=50$.
@@
Simulation results for soft decoding of repetition codes are illustrated in Fig. 7 for the $Z$ and $X$ bases. These plots show the logical error per round $\epsilon_{L}$ and fit the Lambda factor $\Lambda$ from the curve’s slope in logarithmic scale as a function of varying code distances from $d=3$ to $d=51$ for $T=50$ rounds of stabilizer measurements. They compare the soft and hard decoders and include the 68% confidence intervals calculated using the Wilson score interval Wilson (1927). The confidence intervals reflect the varying number of shots for different distances, with higher distances having fewer shots due to subsampling, as indicated by the intervals’ width. Points with fewer than five logical error events appear as dashed lines, suggesting insufficient statistical data to represent the logical error rate.

### v2

@@ I Introduction
Quantum error correction (QEC) is fundamentally recognized as a central strategy for deploying fault-tolerant computations at scale. It promises exponential error suppression, closing the divide between attainable error rates and those required for practical quantum computing. In recent years, significant progress has been achieved in quantum error correction experiments, demonstrating its emerging capabilities[1, 2, 3, 4, 5, 6, 7].
These advancements are particularly evident in implementing the surface code, a leading candidate for fault-tolerant quantum computation. Progress has been marked by an upward path from implementing a distance-three ($d=3$) surface code [1, 2], advancing to a distance-five ($d=5$) surface code [3], and culminating in a distance-seven ($d=7$) surface code that has also shown evidence for improvements in the fidelities of logical entangling gates with increasing code distance [4]. Furthermore, recent experiments claim to have achieved break-even error correction performance, matching or surpassing their physical counterparts. Key milestones include achieving break-even fidelity in the initialization of logical qubits within a color code [4], preparing logical Bell pairs with error rates 800 times lower than their physical level [5], and successfully distilling magic states with fidelity beyond the break-even point [6]. Furthermore, in recent experiments, below surface threshold performance was achieved on a distance-seven ($d=7$) code [7]. These claims highlight quantum error correction’s growing effectiveness and practical relevance.
@@
Pattison et al. developed a method to circumvent this loss by incorporating the full range of analog data, also known as soft information, into the decoding process [8] with a weighting algorithm. Their simulations indicated that using soft information in surface code decoding can significantly raise the code’s threshold compared to decoders that utilize only binary, or “hard”, information. This result reveals a substantial improvement in the exponential scaling of error suppression, going beyond just a constant benefit. The findings underscore the importance of utilizing all available information from the experimental setup, thus optimizing quantum error correction strategies.
In addition to the work by Pattison et al., the benefit of employing soft information has been validated through simulations using neural network decoders [9, 10]. On the hardware front, soft decoding has been implemented in different settings: for improving the readout fidelity on a Si/SiGe two-qubit system [11], and for superconducting qubits using both a distance-three ($d=3$) heavy-hexagon code [12] and, recently, a distance-three ($d=3$) two-dimensional repetition code [13]. However, these hardware experiments were limited to at most $d=3$ codes, which restricted their ability to evaluate and compare thresholds between soft and hard decoding—a significant and anticipated aspect of this research.
The experiments of Ref. [12] involving $T=10$ rounds of stabilizer measurements managed to post-select for leakage errors—transitions to higher energy states detrimental to quantum error correction [14]. This post-selection was feasible because the low number of measurement rounds helped restrict the accumulation of leakage, which typically increases with the duration of the experiment. Furthermore, the experiments of Refs. [13] that used Pattison et al.’s weighting method did not explicitly treat leakage. Consequently, the issue of leakage and its impact on soft decoding with weighting algorithms, also not addressed in the initial studies by Pattison et al. or subsequent simulation efforts, remains a critical area for further investigation.
@@ II.1 Repetition codes
After measuring the stabilizer, one might consider resetting the ancilla to the $|0\rangle$ state to prevent additional error propagation. However, resetting the qubit is a non-trivial task, requiring either waiting for the qubit to decay to the ground state or applying an active reset operation. The former would mean that the code qubits in the system must be idle for as long as the ancillas decay, causing them to decay themselves. The latter would require applying an $X$ gate conditioned on the measurement outcome. However, this would require dynamic operations, which also cause idling times and specific errors. To circumvent these issues, we do not reset the ancilla qubits. Moreover, it has been shown that omitting ancilla resets can have benefits in the decoding of memory experiments [15], further justifying our approach.
@@ II.2 Decoding
Repetition codes are almost cyclic, enabling their decoding through the minimum weight perfect matching (MWPM) algorithm [16, 17]. This decoding method is also extensively used for surface codes, widely recognized for their promising potential in fault-tolerant quantum error correction [1, 2, 3, 4, 7]. Therefore, improved decoding performance achieved with repetition codes likely has implications for the broader application with surface codes.
@@
Error chains, i.e., a set of errors with overlapping syndromes, will induce either one syndrome—if the chain terminates at a boundary—or two syndromes. Therefore the decoding problem is equivalent to finding the shortest path on the graph to pair up all the nontrivial syndromes. The matching decoder solves this problem on a weighted graph, where the edge weights $w=\log((1-p)/p)$ are calculated using the probability $p$ of the corresponding error event. The minimum-weight set of edges found by MWPM then corresponds to the most likely set of errors [16, 17].
We typically evaluate the merit of a quantum code based on its threshold value $p_{th}$, defined as the upper limit of error probability at which the code can still suppress errors. This value originates in the threshold theorem [18], which posits that assuming a constant probability of a fault $p$ across all potential errors, the logical error rate $\epsilon_{L}$ decreases exponentially with respect to the order of fault-tolerance $f$, which is the maximum number of faults that the code can correct [19]. For repetition codes with distance $d$, the order of fault-tolerance is $f=\lfloor d/2\rfloor$ as the code can detect up to $d-1$ errors but only unambiguously correct up to $\lfloor d/2\rfloor$ errors. Consequently, the logical error rate $\epsilon_{L}$ exponentially decreases with the distance of the code as
@@
The code in question does not solely determine the threshold value, as it is also significantly influenced by the decoding method [20]. Typically, researchers simulate the performance of a code and its decoder across varying error rates to estimate the threshold. Moreover, the error parameters in the circuit-level error model do not need to be equal, requiring a threshold parameter for each error mechanism [21]. Determining these thresholds in experiments would require precise manipulation of the noise model and, hence, gates with accurately controllable error rates—a daunting task.
Instead, we employ the error suppression factor $\Lambda$ as a practical metric for comparing the performance of various codes and decoders [22]. We define $\Lambda$ through the scaling of the logical error rate $\epsilon_{L}$ as
@@
When decoding multiple rounds of syndrome measurements, we obtain the total logical error rate $P_{L}$ rather than the previously discussed logical error rate per round $\epsilon_{L}$. Assuming each round incurs a logical error with probability $\epsilon_{L}$, the total logical error rate $P_{L}$ after $T$ rounds is given by [20]:
@@ II.3 Soft information
In superconducting hardware, we perform qubit readouts by coupling a resonator to the qubit. The qubit’s state then induces a dispersive shift in the resonator’s frequency. To detect this shift, we apply a microwave tone to the resonator and measure the reflected signal, which has an altered amplitude and phase according to the qubit’s state. We integrate this reflected signal to extract two key metrics: the in-phase ($I$) and quadrature ($Q$) components and plot them on the IQ plane [23]. This plane has $I$ and $Q$ on its axes and visually represents the qubit’s state through the distribution of $I$ and $Q$ points. Statistically, the $I$ and $Q$ points cluster into two distinct Gaussian distributions on the IQ plane, each representing one of the qubit’s possible states. Fig. 2 shows statistical IQ data obtained from a superconducting qubit. To ascertain the qubit’s state from the reflected signal, we assess the likelihood of a data point belonging to each Gaussian and convert it into a binary outcome. The dispersive shift determines the separation between the two Gaussians, whereas the width of the Gaussian curves is dependent, among other things, on the duration of the integration window. Despite efforts to separate these distributions, a non-zero overlap exists, creating a situation where an IQ data point might fall into an area more likely associated with the incorrect state. Such occurrences lead to a misassignment error, commonly called a soft flip.
@@
The read-out is typically calibrated by preparing a state |$\ket{0}\text{ or }\ket{1}$ and minimizing the probability to measure a |$\ket{1}\text{ or }\ket{0}$ respectively. This doesn’t distinguish between hard and soft flips. However, it is possible to calibrate by minimizing a figure of merit that captures the ambiguity in the distributions of the analog signal [24].
@@
This assumption generally holds for stabilizer measurements after sufficient rounds, as the ancilla qubits will likely be randomly in the $|0\rangle$ or $|1\rangle$ state. However, for the initial rounds and particularly for the final logical code readout, this assumption may not hold. Therefore, incorporating additional information derived from Pauli tracing [25] could improve the model’s accuracy and, consequently, the efficacy of the methods discussed in this chapter.
@@ II.6 Soft decoding in the presence of leakage
Transmon qubits, featured in IBM’s devices, are a leading platform for quantum computation. However, transmons have a finite anharmonicity, allowing transitions to higher energy states. Such transitions can excite a qubit, resulting in leakage errors. These excitations often occur when implementing two-qubit gates [26, 27], as well as measurements [28, 29, 30, 31].
Leakage poses a significant challenge in quantum error correction. Once a qubit leaks, it does not operate as intended, leading to persistent faults over multiple stabilizer measurements. Quantum error correction typically assumes that errors are uncorrelated; however, leakage undermines this assumption, posing a significant challenge for the experimental implementation of quantum error correction protocols. Miao et al. demonstrated the detrimental effect of leakage on decoding performance in superconducting qubits [14], showing a much higher sensitivity of code performance to leakage than Pauli errors. Past experiments involving multiple rounds of quantum error correction have focused on identifying and post-selecting leaked qubits, as seen in the studies conducted by [20, 1].
Soft information decoding experiments have also selectively removed samples with detected leaked qubits [13, 12]. Notably, these studies involved no more than $T=16$ rounds of measurements. In contrast, our experiments extend to $T=100$ rounds, significantly increasing the likelihood of leakage and thus reducing the yield to nearly zero if post-selection were applied.
Several strategies have been developed to mitigate leakage in quantum systems, such as implementing leakage-reducing reset protocols [20, 32], applying optimal control techniques to gate operations [33], and integrating leakage reduction units directly into the circuit design [34, 35]. In line with our decision to use IBM Quantum’s deployed device IBM Sherbrooke, which is not optimized for quantum error correction, we explored the extreme case where no leakage mitigation is applied at the hardware level. Instead of focusing on implementing sophisticated leakage-reduction protocols, we investigated how effectively leakage can be managed solely through classical decoding techniques. As Fowler et al. [35] showed, a threshold still exists under leakage, although significantly lower, making this a compelling test of soft decoding’s robustness under realistic, unmitigated conditions.
Leakage manifests as a separate Gaussian distribution in the IQ plane, typically positioned between the Gaussians representing the computational states [12]. This separation allows us to detect leakage using the IQ data already available, enabling the classification of leaked qubits based on their IQ position.
@@
Second, as described in [12], the Gaussian distribution of the leaked state has a variance similar to the computational states, ideally calling for a linear decision boundary between leaked and computational states. Our approach, based on outlier detection, results in a circular decision boundary, which tends to underestimate the number of leaked states, as shown in Fig. 6. Addressing this limitation would ideally involve including a distinct preparation of leaked states in the calibration data and performing a three-state discrimination process, as described in [12].
@@ III.1 Simulations
To explore the advantages of soft decoding repetition codes on superconducting hardware, we initially performed simulations that mimic the behavior of realistic quantum hardware to assess the effectiveness of soft decoding. After these preliminary simulations, we conducted experiments on IBM Quantum’s deployed device IBM Sherbrooke, a 127-qubit Eagle device not specifically tuned for quantum error correction tasks like repetition codes. Rather than mitigate leakage and other measurement errors through fine-tuning of the measurement process, as would be done in a full demonstration of an error correcting code, we instead simply use the measurements as provided “out-of-the-box” through Qiskit [36]. This allows us to investigate how much the leakage can be mitigated purely through the decoding. Using this general-purpose approach, we did not aim to maximize QEC performance but to evaluate how soft decoding behaves under practical conditions comparing the outcomes with those of the simulations.
@@
While this setup with $T=50$ does not approach the condition of $T\gg d$, it aligns with the conventional experimental practice where $d$ is set equal to $T$, ensuring that the “time distance” equals the code distance. Ref. [20] showed that time boundary effects in repetition codes do not affect the logical error rate per round for $T\geq 10$, supporting our decision to set $T=50$.
@@
Simulation results for soft decoding of repetition codes are illustrated in Fig. 7 for the $Z$ and $X$ bases. These plots show the logical error per round $\epsilon_{L}$ and fit the Lambda factor $\Lambda$ from the curve’s slope in logarithmic scale as a function of varying code distances from $d=3$ to $d=51$ for $T=50$ rounds of stabilizer measurements. They compare the soft and hard decoders and include the 68% confidence intervals calculated using the Wilson score interval [37]. The confidence intervals reflect the varying number of shots for different distances, with higher distances having fewer shots due to subsampling, as indicated by the intervals’ width. Points with fewer than five logical error events appear as dashed lines, suggesting insufficient statistical data to represent the logical error rate.